f(x)=x^2+2x-4 Give the vertex, axis, domain, and range.
Answer by lwsshak3(11628)

f(x)=x^2+2x-4 Give the vertex, axis, domain, and range.
***
f(x)=x^2+2x-4
complete the square:
f(x)=(x^2+2x+1)-1-4
f(x)=(x+1)^2-5
This is an equation of a parabola that opens upward.
Its standard (vertex) form of equation: y=(x-h)^2+k, (h,k)=coordinates of the vertex
For given parabola:
vertex: (-1,-5)
axis of symmetry: x=-1
Domain: (-∞,∞) or all real numbers
Range: (-5,∞)